find the larger of two consecutive integers if the second is 13 less than twice the smaller
Answer by Mathtut(3670)

lets call the smaller x and the larger x+2 now x+2(larger)=2x-13 or x+2=2x-13 so x=15 which means that x+2=17
so the larger integer = 17